OpenStudy (anonymous):

Sara is now 3 times older than Kristin. Four years ago, Sara was 5 times as old as Kristin was then. How old is Sara now?

OpenStudy (anonymous):

S and K be the current ages of Sara and Kristin, respectively. Solve the following for S and K: { S = 3 K, (S - 4) = 5 ( K - 4) } { S = 24 years old, K = 8 years old }

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Your welcome.